OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. -7x-5y=1.5 -5x-4y=4.5

OpenStudy (ash2326):

1st......-7x-5y=1.5 2nd.....-5x-4y=4.5 multiply 1st by 5 3rd......-35x-25y=7.5 multiply second by 7 4th......-35x-28y=31.5 subtract 3rd from 4th -35x-28y-(-35x-25y)=31.5-7.5 -3y=24 y=-8 now substitute this in 1st -7x+40=1.5 -7x=-38.5 x=5.5 so x=5.5 and y=-8

OpenStudy (anonymous):

Thanks so much